OpenStudy (anonymous):

find (dw/dy)x,t if w=x^2+y-z+sint and t=x+y

OpenStudy (kainui):

Show me your best attempt and I will help you.

OpenStudy (anonymous):

Find \[ \left( \delta w/\delta y \right)_{x,t} if w=x ^{2}+y-z+\sin t\] and t=x+y

OpenStudy (kainui):

Is z a function of w, x, y, or t?

OpenStudy (anonymous):

z is a function of w